A LOOK AT A SERIES OF ALKYL AND PERFLUOROALKYL BROMIDES AND CHLORIDES

\underline{BRITTANY E. LONG}, STEPHEN A. COOKE, Department of Chemistry, The University of North Texas, 1155 Union Circle, #305070, Denton, TX 76203-5017, U.S.A.; GARRY S. GRUBBS II, Department of Chemistry, Wesleyan University, Hall-Atwater Laboratories, 52 Lawn Ave., Middletown, CT 06459-0180, U.S.A.

The pure rotational spectrum for bromoperfluoroethane between 8.0 and 14.0 GHz and chloroperfluoroethane between 8.0 and 16.0 GHz has been measured on a chirped pulse Fourier transform microwave spectrometer for the first time. A total of 839 transitions for the bromoperfluoroethane, which includes the ^{79}Br, ^{81}Br parent isotopologues and the four 13C’s, have been assigned quantum numbers. 496 transitions were observed for chloroperfluoroethane, which includes the 35Cl and ^{37}Cl species. Only the trans conformers were observed for which the rotational constants are reported. Nuclear electric quadrapole coupling constants have been determined and reported. Also, two dipole forbidden/quadrapole allowed ΔJ = 2 transitions were observed in only the bromoperfluoroethane spectra. No forbidden transitions were observed in the chloroperfluoroethane.